:A bubble originating at the bottom of a lake rises to the surface within 10.0 seconds with an acceleration of 10.0 meters/second2. What is the depth of the lake?Select one of the options below as your answer:100 meters 200 meters 300 meters 400 meters 500 meters
1
Expert's answer
2012-06-19T08:37:47-0400
Let the depth of the lake is d.
d = a * t^2 / 2 = 10 * (10^2) / 2 = 1000 / 2
= 500
Answer: 500 meters.